Using the following production possibilities table for Cars and Guided Missiles
answer the questions below:
Type of
Product
A B C D E
Cars
(In millions)
0 2 4 6 8
Guided
Missiles
(In thousands)
30 27 21 12 0
(a) Graph the data in the table. Are there constant or increasing costs for the
production of missiles? (2 marks)
(b) If the economy is currently at the point C:
 What is the cost of one million more cars? (2 mark)
 What is the cost of one thousand more guided missiles? (2
mark)
(c) Label a point G inside the curve of the graph for questions (a). What doesit indicate? (2 marks)
(d) Label a point H that lies outside the production possibilities curve of
question (2 marks)
 What does this point indicate?
Answer #1

A)

cars (in millions) o vino KPPF Ot 12 21 27 30 & Gewided missiles (in thousants)

Suppose the economy is at point E, producing 8 millions car and 0 thousands guided missiles. Now if the economy moves from point E to point D, it produces 6 millions cars and 12 thousands guided missiles. It means opportunity cost of producing 1 thousand more guided missiles is [(8 - 6)/(12 - 0)] = 0.17 millions of cars. Now if the economy moves from point D to point C, it will produce 4 millions cars and 21 thousands guided missiles. Now, the opportunity cost of producing 1 thousand more guided missiles is [(6 - 4)/(21 - 12)] = 0.22 millions of cars.

It means opportunity cost of production of missiles is increasing.

B) If the economy is at point C, it is producing 4 millions cars and 21 thousands of guided missiles. Opportunity cost of one million more cars is [(21 - 12)/(6 - 4)] = 4.5 thousands of guided missiles.

Opportunity cost of one thousand more guided missiles is [(4 - 2)/(27 - 21)] = 0.33 million of cars.

(C) - (d)

(C) - (J) carsim milliams) H G wided missiles 30 im thousand) point G represents a feasible combination of both goods but it